Rumsfeld Corporation leased a machine on December 31, 2018, for a three-year period. The lease agreement calls for annual payments in the amount of $17,000 on December 31 of each year beginning on December 31, 2018. Rumsfeld has the option to purchase the machine on December 31, 2021, for $21,000 when its fair value is expected to be $31,000. The machine's estimated useful life is expected to be six years with no residual value. The appropriate interest rate for this lease is 10%.
  

n/i PV of $1 PV, ordinary annuity PV, annuity due
1 period, 10% 0.90909 0.90909 1.00000
2 periods, 10% 0.82645 1.73554 1.90909
3 periods, 10% 0.75131 2.48685 2.73554


Required:
1. Calculate the amount to be recorded as a right-of-use asset and the associated lease liability.
2. Prepare an amortization schedule for this lease.
Answer #1
1) PV of $1, n=3, i=10% = .75131
PV of an annuity due of $1, n=3, i=10% = 2.73554
PV of payments $17,000 x 2.73554 = $ 46,504.18
PV of Purchase Option $21,000 x.75131 = $ 15,777.51
Right of use assets / Associated Lease Liability $        62,282
A B C = A-B D = C x 10% E = C+D
2) Time Initial value Lease payment Balance after payment Interest Expense Balance after adding interest
Dec 31, 2018 $        62,282 17000 $              45,282 4528.169 $                  49,810
Dec 31, 2019 $        49,810 17000 $              32,810 3280.9859 $                  36,091
Dec 31, 2020 $        36,091 17000 $              19,091 1909.08449 $                  21,000
Hint for first row:
62282 = as calculated above
17000 = given
45282 = 62282-17000
4528 = 45282 x 10%
49810 = 45282+4528